OpenStudy (boi.robben):

The curve y = e−x for 0  x  2 is rotated around the x-axis to produce a 3-dimensional solid whose volume is V =  2 (1−e−4). The centre of mass lies somewhere on the x-axis - where?

OpenStudy (amistre64):

make sure all the parts posted correctly in the question ....

OpenStudy (boi.robben):

yea.. they are correct

OpenStudy (amistre64):

y = e−x for 0 x 2 notationally this is unreadable ...

OpenStudy (boi.robben):

wait a sec ... will check it again!!

OpenStudy (boi.robben):

The curve y = e^−x for 0 <x< 2 is rotated around the x-axis to produce a 3-dimensional solid whose volume is V =pie/ 2 (1−e^−4). The centre of mass lies somewhere on the x-axis - where?

OpenStudy (amistre64):

center of mass would most likely be situated when we have half the volume

OpenStudy (amistre64):

id prolly integrate it from 0 to b; and equate that to half the volume that is stated to solve for b

OpenStudy (amistre64):

\[\int_{a}^{n}f(x)~dx=\frac{V}{2}=F(n)-F(a)\] \[\frac{V}{2}+F(a)=F(n)\] and invert for n

OpenStudy (amistre64):

there is also some sort of moment calculations that are a bit foggy ... since it might be hard to integrate e^(-2x) in general

OpenStudy (boi.robben):

i got this e^-x ?? :/

OpenStudy (boi.robben):

please solve it completely!!!

OpenStudy (amistre64):

i was thinking e^(x^2) ... this is simpler \[-\frac{1}{2}\pi e^{-2(n)}+\frac{1}{2}\pi e^{-2(0)}=\frac12*\frac{\pi}{2}(1-e^{-4})\] \[-\pi e^{-2(n)}+\pi e^{-2(0)}=\frac{\pi}{2}(1-e^{-4})\] \[-e^{-2(n)}+ 1=\frac{1}{2}(1-e^{-4})\] \[e^{-2(n)}=1-\frac{1}{2}(1-e^{-4})\] \[-2n=ln(1-\frac{1}{2}(1-e^{-4}))\] \[n=-\frac12ln(1-\frac{1}{2}(1-e^{-4}))\]

OpenStudy (amistre64):

chk the algebra ... but that is what i get too